# Flexibility Pricing in Distribution Systems: A Direct Method Aligned with Flexibility Models

### v1

@@ Abstract
Effective quantification of costs and values of distributed energy resources (DERs) can enhance the power system’s utilization of their flexibility. This paper studies the flexibility pricing of DERs in distribution systems. We first propose a flexibility cost formulation based on adjustment ranges in both power and accumulated energy consumption. Compared with traditional power range-based approaches, the proposed method can quantify the potential inconvenience caused to DER users by flexibility activation and directly capture the time-dependent characteristics of DER’s flexibility regions. We then propose a flexibility cost formulation aligned with the aggregated flexibility model and design a DSO flexibility market model to activate the aggregated flexibility in the distribution system to participate in the transmission system operation for energy arbitrage and ancillary services provision. Furthermore, we propose the concept of marginal flexibility price for the settlement of the DSO flexibility market, ensuring incentive compatibility, revenue adequacy, and a non-profit distribution system operator. Numerical tests on the IEEE 33-node distribution system verify the proposed methods.
@@
Flexibility pricing, distributed energy resources, flexibility market, distribution systems, marginal prices.
@@ I Introduction
Electricity supply and demand in traditional power systems have been balanced by adjusting the outputs of generators and centralized energy storage systems to track load variations. However, in response to the global energy crisis, power systems worldwide are transforming to integrate more and more renewable energies, such as wind power and photovoltaic (PV), which are highly volatile and thereby increase the demand for balancing services. In this context, power system operators have started to pay more attention to the flexibility offered by distributed energy resources (DERs) on the demand side, e.g., electric vehicles (EVs), battery energy storage systems (BESS), and heat pumps (HPs). Characterized by their flexible operational capabilities, DERs possess the potential to offer substantial flexibility to power systems and to mitigate the variability from renewables. Consequently, activating the flexibility of demand-side DERs in power systems operation has become a widespread goal [1].
The flexibility of demand-side DERs can be utilized both to provide balancing services at the transmission system level and to support voltage control and congestion management at the distribution system level [2]. This paper mainly focuses on the former application, yet the proposed methodologies can also be intuitively adapted to the latter scenario.
Due to the small individual capacity and large population of DERs, it is impractical for transmission system operators (TSOs) to consider each DER when computing the dispatch strategy. DERs need to be aggregated to a certain capacity before participating in the transmission system operation, such as for energy arbitrage and ancillary services provision, leading to the emergence of aggregators who serve as intermediaries between DER users and the TSO [3]. On the other hand, since DERs are physically located within distribution systems, the dispatch of DERs can affect the security and power quality of distribution systems. Thus, distribution system operators (DSOs) should also play an important role in flexibility activation. In this paper, we consider the framework where the DSO coordinates a flexibility market to activate the flexibility of DERs (with intermediary aggregators) to participate in the transmission system operation, as depicted in Fig. 1.
@@
In Fig. 1, the flexibility of individual DERs is first collected by their aggregators, who compensate DER users for activating their flexibility through contracts. Aggregators then calculate their aggregated flexibility and associated costs to bid in the distribution system’s flexibility market. The flexibility market is cleared by the DSO based on the price signals from the TSO and the bids from aggregators. The DSO then allocates the obtained revenue from the transmission level to aggregators. This framework, or similar ones, can be found in many existing studies [4, 5, 6, 7]. Ensuring that the DSO is non-profit in the flexibility market is essential for the framework to attract aggregators to participate and operate sustainably. In another flexibility activation framework, the DSO calculates security restrictions for each aggregator, and aggregators directly participate in the transmission system operation under these restrictions [8, 9]. This framework, however, leads to a potential conservativeness of each aggregator’s flexibility due to insufficient coordination among aggregators. Therefore, this paper focuses on the market framework shown in Fig. 1 that allows the DSO to coordinate aggregators to provide a larger total flexibility to the transmission system. Within this framework, the primary issue is how to formulate the aggregated flexibility and its associated price.
Numerous studies have focused on the modeling of aggregated flexibility, proposing models such as the power boundary model [10], power-energy boundary model [11, 12], vertex-based polytope [13], homothetic polytope [14], and $k$-order feasible region [15]. Yet, these studies do not discuss how to quantify the costs or values of the feasible regions determined by these aggregated flexibility models, which is essential to activate flexibility. Although the power boundary model offers a straightforward quantification method by defining prices for the adjustable power range in each time slot, just like conventional generators, it faces challenges in capturing the time-coupled characteristics of storage or storage-like DERs, e.g., BESSs, EVs, and HPs. Within the research focusing on DSO flexibility markets, defining and formulating flexibility prices is always a fundamental part. Existing research in this field predominantly formulates flexibility prices based on the power adjustment range relative to the reference or baseline power in each time slot [4, 16, 5, 17, 18, 19].
@@
The advantages of flexibility pricing based on power adjustment ranges are its intuitive calculation and uniformity with generators. However, as mentioned before, this pricing scheme is insufficient to quantify the time-coupled flexibility offered by storage-like DERs. Besides, the potential inconvenience caused to DER users by flexibility activation, which is the fundamental reason for flexibility costs, cannot be adequately quantified by the power range alone (detailed discussions on this fact will be provided in Section II). In [9], the opportunity cost is used to evaluate the cost of EVs’ mid-term flexibility, but this method cannot be directly generalized to various heterogeneous DERs, limiting its application in a unified DSO flexibility market with multiple types of DERs.
@@
In summary, there is a lack of discussion on the pricing method of the time-coupled flexibility and the associated flexibility market. Filling this gap enables a more reasonable evaluation of the costs and values of DERs’ flexibility, fostering a more effective DSO flexibility market and enhancing the willingness of DER users to offer their flexibility to contribute to power systems operation. Therefore, this paper proposes a direct flexibility pricing method aligned with the time-coupled flexibility models. Moreover, we design a DSO flexibility market model to activate the flexibility of DERs to participate in the transmission system operation for energy arbitrage and ancillary services provision. The contributions are threefold:
@@
  • From the perspective of DER users, we propose to compute the flexibility cost based on activated adjustment ranges in both power and accumulated energy consumption. This method is in line with the power-energy boundary models of individual DERs’ flexibility and is able to quantify the potential inconvenience caused to users by flexibility activation, which is more suitable for typical DERs such as EVs, BESSs, and HPs compared to the traditional power range-based approach.
@@
  • From the aggregators’ standpoint, we propose to compute their flexibility costs directly according to the activated area of the aggregated flexibility model. Unlike the traditional power range-based formulations, the proposed approach aligns with the aggregated flexibility model with time coupling, thus allowing aggregators to define their cost coefficients directly based on the feasible regions determined by their flexibility models.
  • From the DSO’s perspective, we propose a market-clearing model for activating the flexibility of all aggregators in the distribution system to participate in the transmission system operation for energy arbitrage and reserve services provision. Based on this model, we propose the concept of marginal flexibility price for distributing revenues earned from the transmission level to each aggregator, ensuring incentive compatibility, revenue adequacy, and a non-profit DSO.
@@
II Flexibility Cost Formulation of DERs based on Both Power and Energy Adjustment Ranges
This section first presents the proposed unified flexibility cost formulation based on the adjustment ranges in both power and accumulated energy consumption. We then discuss how to specify the cost coefficients in this formulation according to the distinct operational characteristics of typical DERs.
II-A Unified Formulation for Flexibility Costs
Flexibility refers to the capability of a device to adjust its operation. From the power system’s perspective, the flexibility cost of a device should be computed by the adjustment range it provides, similar to the capacity costs for generators providing ancillary services. [^1] In ancillary services, mileage costs may also be counted, but in a post-operational fashion. Here, the primary purpose of formulating the flexibility cost is to support pre-operational decision-making. In this context, mileage costs are typically estimated by multiplying capacity costs by a proportion based on historical regulation data [4]. Therefore, we focus on capacity costs. The proposed flexibility cost formulation is based on adjustment ranges in both power and accumulated energy, which aligns with the classic power-energy flexibility model of individual DERs [11]. Besides, this proposed flexibility cost formulation can quantify the potential inconvenience caused to DER users upon flexibility activation.
We discretize the time horizon into $T$ slots, each of a length $\Delta_{T}$, indexed by $t$, and denote the whole time horizon as $[T]\triangleq\{1,2,...,T\}$. The flexibility cost of one DER is:
@@ II-A Unified Formulation for Flexibility Costs
where $\Delta\hat{\mathbf{p}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\mathbf{p}}\in\mathbb{R}^{T}_{+}$ denote the vectors of activated upward and downward adjustment ranges of the power profile $\mathbf{p}\in\mathbb{R}^{T}$ relative to the baseline profile $\mathbf{p}^{{\text{base}}}\in\mathbb{R}^{T}$, respectively; $\Delta\hat{\mathbf{e}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\mathbf{e}}\in\mathbb{R}^{T}_{+}$ represent the activated upward and downward adjustment ranges of the accumulated energy consumption trajectory $\mathbf{e}\in\mathbb{R}^{T}$ relative to the baseline energy trajectory $\mathbf{e}^{{\text{base}}}\in\mathbb{R}^{T}$, respectively; $\hat{\mathbf{c}}_{\text{p}}\in\mathbb{R}^{T}$, $\check{\mathbf{c}}_{\text{p}}\in\mathbb{R}^{T}$, $\hat{\mathbf{c}}_{\text{e}}\in\mathbb{R}^{T}$, and $\check{\mathbf{c}}_{\text{e}}\in\mathbb{R}^{T}$ denote the vectors of cost coefficients for upward and downward adjustment ranges in power and energy, depending on the specific DERs. Variables $\mathbf{p}$, $\Delta\hat{\mathbf{p}}$, $\Delta\check{\mathbf{p}}$, $\mathbf{e}$, $\Delta\hat{\mathbf{e}}$, and $\Delta\check{\mathbf{e}}$ should fulfill the following constraints:
@@
where $\overline{\mathbf{p}}\in\mathbb{R}^{T}$, $\underline{\mathbf{p}}\in\mathbb{R}^{T}$, $\overline{\mathbf{e}}\in\mathbb{R}^{T}$, and $\underline{\mathbf{e}}\in\mathbb{R}^{T}$ denotes the DER’s physical upper and lower bounds of power and accumulated energy. Constraints (2) and (3) restrict the power and energy adjustment ranges within their physical bounds. Equation (4) gives the relationship between power and energy.
The constraints in (2)-(4) actually constitute the power-energy boundary model that describes the flexibility of typical DERs, with bound parameters determined by the operational characteristics of different DERs, as detailed in [11]. The formulation (1) aligns with the flexibility model (2)-(4) since it calculates the flexibility cost based on the activated adjustment ranges in both power and accumulated energy consumption. By properly setting the cost coefficient vectors $\hat{\mathbf{c}}_{\text{p}}$, $\check{\mathbf{c}}_{\text{p}}$, $\hat{\mathbf{c}}_{\text{e}}$, and $\check{\mathbf{c}}_{\text{e}}$, the flexibility cost (1) can reflect the potential inconvenience caused to DER users upon flexibility activation, which will be discussed in the next subsection.
@@ II-B Specification of the Cost Coefficients for Typical DERs
For resources whose costs are directly quantified by power adjustment ranges, like distributed PV and curtailable loads, specifying the cost coefficients is straightforward: set $\hat{\mathbf{c}}_{\text{e}}$ and $\check{\mathbf{c}}_{\text{e}}$ to zero and specify $\hat{\mathbf{c}}_{\text{p}}$ and $\check{\mathbf{c}}_{\text{p}}$ identically as the traditional power range-based formulations. Hence, the rest of this subsection mainly discusses how to specify the cost coefficients for typical storage-like DERs, including EVs, BESS, and HPs, based on the potential inconvenience caused to users of these DERs upon flexibility activation.
@@ II-B1 Electric Vehicles
The flexibility of EVs lies in the charging process during their plug-in periods. Since the primary goal of EV charging is to charge the battery to the expected energy level by departure, the inconvenience caused to EV users by activating this flexibility is reflected in the potentially unsatisfied charging demand at the departure time.
Fig. 2 illustrates a typical EV’s adjustment ranges in the accumulated energy consumption and its flexibility cost coefficients. In this example, the EV has an expected value for the energy to be charged by departure but also allows it to be adjustable within certain upper and lower bounds. The upper bound typically corresponds to the battery’s capacity limit, while the lower bound represents the minimum energy level acceptable to the EV user. The baseline energy trajectory is the energy trajectory where charging starts immediately at the rated power upon plug-in until the expected energy level is reached. The area within the upward and downward energy adjustment ranges represents the activated flexibility in accumulated energy consumption, which is where the actual energy trajectory can be adjusted. However, only the energy adjustment at the departure time can cause inconvenience to the EV user. More precisely, it is the downward adjustment that matters, as the upward adjustment, representing charging more than the expected energy, typically does not lead to an inconvenience in the EV’s usage. Thus, in the flexibility cost coefficients, only the coefficient for the downward energy adjustment range $\check{\mathbf{c}}_{\text{e}}$ at the departure time is non-zero, while $\check{\mathbf{c}}_{\text{e}}$ at other times, $\hat{\mathbf{c}}_{\text{e}}$, $\check{\mathbf{c}}_{\text{p}}$ and $\hat{\mathbf{c}}_{\text{p}}$ are all zero. [^2] The increased energy cost due to the overcharging indicated by the upward adjustments in the energy trajectory is not considered within the scope of flexibility costs. Instead, these costs will be incorporated into the energy cost component in the objective function of the flexibility activation program.
@@
In summary, based on the flexibility cost model (1), the aggregator can establish a contract with EV users that specifies compensation for any unsatisfied charging demands at the departure time. This contract directly correlates to the value of $\check{\mathbf{c}}_{\text{e}}$. Consequently, EV users can intuitively understand the potential inconvenience to their EV usage due to flexibility activation and evaluate their expected benefits. This level of clarity and directness is not achievable with traditional cost quantification methods based on power adjustment ranges, as power adjustments during charging are invisible to EV users.

### v4

# Quantifying and Optimizing the Time-Coupled Flexibilities at the Distribution-Level for TSO-DSO Coordination

@@ Abstract
The flexibilities provided by the distributed energy resources (DERs) in distribution systems enable the coordination of transmission system operator (TSO) and distribution system operators (DSOs). At the distribution level, the flexibilities should be optimized for participation in the transmission system operation. This paper first proposes a flexibility quantification method that quantifies the costs of providing flexibilities and their values to the DSO in the TSO-DSO coordination. Compared with traditional power-range-based quantification approaches that are mainly suitable for generators, the proposed method can directly capture the time-coupling characteristics of DERs’ individual and aggregated flexibility regions. Based on the quantification method, we further propose a DSO optimization model to activate the flexibilities from DER aggregators in the distribution system for energy arbitrage and ancillary services provision in the transmission system, along with a revenue allocation strategy that ensures a non-profit DSO. Numerical tests on the IEEE test system verify the proposed methods.
@@
Flexibility quantification, distributed energy resources, distribution systems, flexibility activation, TSO-DSO coordination.
@@ I Introduction
Electricity supply and demand in traditional power systems have been balanced by adjusting the outputs of generators and centralized energy storage systems to track load variations. However, in response to the global energy crisis, power systems worldwide are transforming to integrate more and more renewable energies, such as wind power and photovoltaic (PV), which are highly volatile and thereby increase the demand for balancing services. In this context, power system operators have started to pay more attention to the flexibility offered by distributed energy resources (DERs) on the demand side, e.g., electric vehicles (EVs), battery energy storage systems (BESS), and heat pumps (HPs). Characterized by their flexible operational capabilities, DERs possess the potential to offer substantial flexibility to power systems and to mitigate the variability from renewables. Consequently, activating the flexibility of demand-side DERs in power system operations has become a widespread goal [1].
Due to the small individual capacity and large population, DERs need to be aggregated to a certain capacity for better management and computation from the power system’s perspective, leading to the emergence of third-party aggregators who serve as intermediaries between DER users and the power system operators-the transmission system operator (TSO) or the distribution system operator (DSO) [2]. Aggregating a large number of DERs can provide flexibility for the transmission system operation, such as participating in energy arbitrage and providing ancillary services. However, since DERs are physically located within distribution systems, adjustments in the power of numerous DERs can affect the distribution system’s security, power quality, and operational efficiency. Thus, the DSO also needs to be involved in aggregating the DERs’ flexibilities to participate in transmission system operations, i.e., TSO-DSO coordination.
This paper considers the TSO-DSO coordination framework as depicted in Fig. 1. In this framework, the flexibility of individual DERs is first collected by the aggregators, who compensate the DER owners for activating their flexibility through contracts. Aggregators then calculate their aggregated flexibility and its associated costs to report to the DSO. The DSO coordinates the aggregators’ flexibility in calculating the total power profile and ancillary service capacities based on the price signals from the TSO. The obtained revenues from the TSO are finally allocated to aggregators. This framework, or similar ones, can be found in many existing studies [3, 4, 5, 6]. Ensuring that the DSO is non-profit in this process is essential for the framework to attract DER aggregators to participate and operate sustainably.
@@
In another TSO-DSO coordination framework, the DSO calculates security restrictions for each aggregator, and aggregators directly participate in the transmission system operation under these restrictions [7, 8]. This framework, however, leads to a potential conservativeness of each aggregator’s flexibility due to insufficient coordination among aggregators. Therefore, this paper focuses on the framework shown in Fig. 1 that allows the DSO to coordinate aggregators to provide a larger total flexibility to the transmission system. Within this framework, the primary issue is how to model the aggregated flexibility and quantify its cost and value for the system. In subsequent text, unless specified, the term “quantification” refers to quantifying both the cost of flexibility and its value for the system.
Numerous studies have focused on the modeling of aggregated flexibility, proposing models such as the power boundary model [9], power-energy boundary model [10, 11], vertex-based polytope [12], ellipsoid [13], generalized battery model [14], and homothetic polytope [15]. Our previous works [16, 17] also proposed efficient modeling methods for flexibility aggregation. Nevertheless, these studies did not discuss how to quantify the costs or values of the feasible regions determined by these aggregated flexibility models, which is important for the DSO’s optimal activation of flexibilities.
Within the field of DSO flexibility markets, quantifying the costs and values of flexibility is always a fundamental part. However, existing studies did not discuss this aspect in detail; instead, they have essentially adopted quantification methods based on the power adjustment ranges in each time slot[18, 3, 4, 19, 20]. For example, reference [18] proposes to discretize the power adjustment range in each time slot into standardized contract capacities and to define different prices for each segment. References [3] and [4] incorporate reserve capacities within the time-coupled flexibility model of aggregators and define their flexibility costs as reserve costs. This quantification method is still essentially based on the power adjustment ranges in each time slot because defining reserves implicitly calculates an approximated power range model (time-decoupled) of the original flexibility model.
@@
The advantages of using the power adjustment range in each time slot for flexibility quantification are its intuitive calculation and uniformity with generators. However, this method is insufficient for the quantification of the time-coupled flexibility offered by storage-like DERs, e.g., BESSs, EVs, and HPs. In [8], the opportunity cost is used to quantify the cost of EVs’ mid-term flexibility, but this method cannot be directly generalized to various heterogeneous DERs, limiting its application in a unified flexibility activation framework.
@@
In summary, there is a lack of discussion on the quantification method of the cost and value of the time-coupled flexibility model. Filling this gap enhances the willingness of DER users to offer their flexibility to contribute to power system operations. Therefore, this paper proposes a unified quantification method aligned with the time-coupled flexibility models. The term “unified” here refers to the fact that the proposed quantification method maintains a unified form for the individual and aggregated flexibility of typical DERs. Moreover, we design a DSO optimization model to activate the flexibility of DERs to participate in the transmission system operation for energy arbitrage and ancillary services provision, along with a revenue allocation strategy. The contributions of this paper are fourfold:
@@
  • From the perspective of DER users, we propose to uniformly compute the flexibility cost based on activated adjustment ranges in both power and accumulated energy consumption. This method is in line with the time-coupled power-energy boundary models of individual DERs’ flexibility, which is more suitable for storage-like DERs such as EVs, BESSs, and HPs compared to the traditional power-range-based approach.
@@
  • From the aggregators’ standpoint, we propose to compute their flexibility costs directly according to the activated area of the aggregated flexibility model. Unlike the traditional power-range-based formulations, the proposed approach aligns with the aggregated flexibility model with time coupling, thus allowing aggregators to define their cost coefficients directly based on the feasible regions determined by their flexibility models.
  • Based on the flexibility cost quantification method, we propose a DSO optimization model for coordinating the flexibility of all aggregators in the distribution system to participate in the transmission system operation for energy arbitrage and ancillary services provision.
  • From the DSO optimization model, we derive the concept of marginal flexibility prices for quantifying the marginal contributions of each aggregator’s flexibility to the total revenue of the DSO. The marginal flexibility prices are then used to allocate the revenues earned from the transmission system to each aggregator, ensuring a non-profit DSO in the flexibility activation process.
@@
II Quantification of Flexibility Costs for Individual DERs
This section first presents the flexibility cost quantification method for individual DERs based on adjustment ranges in both power and accumulated energy consumption. Then, we discuss how to specify the cost coefficients in this method according to the distinct characteristics of various DERs.
II-A Flexibility Cost Quantification based on Both Power and Energy Adjustment Ranges
When DERs provide flexibilities to the power system, their power may be adjusted within the offered flexibility region. Power adjustments can lead to equipment degradation or inconvenience for users, which constitute the primary source of the DERs’ flexibility costs. Flexibility costs are considered in the pre-operational decision-making stage, which reflects the expected adjustment costs in the subsequent real-time operation stage, so they should be quantified based on the adjustment capability defined by the flexibility region.
Conceptually, flexibility costs are similar to the capacity costs for generators providing ancillary services. [^1] In ancillary services, mileage costs may also be counted, but in a post-operational fashion. Here, the primary purpose of formulating the flexibility cost is to support pre-operational decision-making. In this context, mileage costs are typically estimated by multiplying capacity costs by a proportion based on historical regulation data [3]. Therefore, we focus on capacity costs. However, since many DERs’ flexibility region exhibits time-coupled characteristics, their flexibility costs can not be adequately captured based on the power adjustment range of each time slot, which is, though, suitable for conventional generators. Hence, we propose to quantify the individual DERs’ flexibility costs based on adjustment ranges in both power and accumulated energy, which aligns with the classic power-energy flexibility model (with time-coupling) of individual DERs [10].
First, the operational time horizon is discretized into $T$ slots, denoted as $[T]\triangleq\{1,2,...,T\}$. Each time slot has a length of $\Delta_{T}$ and is indexed by $t$. The flexibility cost of one DER is:
@@ II-A Flexibility Cost Quantification based on Both Power and Energy Adjustment Ranges
where $\Delta\hat{\mathbf{p}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\mathbf{p}}\in\mathbb{R}^{T}_{+}$ denote the vectors of activated upward and downward adjustment ranges of the power profile $\mathbf{p}\in\mathbb{R}^{T}$ relative to the baseline profile $\mathbf{p}^{{\text{base}}}\in\mathbb{R}^{T}$, respectively; $\Delta\hat{\mathbf{e}}\in\mathbb{R}^{T}_{+}$ and $\Delta\check{\mathbf{e}}\in\mathbb{R}^{T}_{+}$ represent the activated [^2] Because not all the flexibilities offered by the DER are valuable for the system, only a portion of the flexibility will be actually utilized in the power system operations. This paper calls this portion of the DER’s flexibility ”activated.” upward and downward adjustment ranges of the accumulated energy consumption trajectory $\mathbf{e}\in\mathbb{R}^{T}$ relative to the baseline energy trajectory $\mathbf{e}^{{\text{base}}}\in\mathbb{R}^{T}$, respectively; $\hat{\mathbf{c}}_{\text{p}}\in\mathbb{R}^{T}$, $\check{\mathbf{c}}_{\text{p}}\in\mathbb{R}^{T}$, $\hat{\mathbf{c}}_{\text{e}}\in\mathbb{R}^{T}$, and $\check{\mathbf{c}}_{\text{e}}\in\mathbb{R}^{T}$ denote the vectors of cost coefficients for upward and downward adjustment ranges in power and energy, depending on the specific DERs. Variables $\mathbf{p}$, $\Delta\hat{\mathbf{p}}$, $\Delta\check{\mathbf{p}}$, $\mathbf{e}$, $\Delta\hat{\mathbf{e}}$, and $\Delta\check{\mathbf{e}}$ should fulfill the following constraints:
@@
where $\overline{\mathbf{p}}\in\mathbb{R}^{T}$, $\underline{\mathbf{p}}\in\mathbb{R}^{T}$, $\overline{\mathbf{e}}\in\mathbb{R}^{T}$, and $\underline{\mathbf{e}}\in\mathbb{R}^{T}$ denotes the DER’s physical upper and lower bounds of power and accumulated energy. Constraints (2) and (3) restrict the DER’s power and energy profile within their activated adjustment ranges and the activated power and energy adjustment ranges within their physical bounds. Equation (4) gives the relationship between power and energy.
The constraints in (2)-(4) actually constitute the power-energy boundary model that describes the flexibility of typical DERs such as PVs, EVs, BESSs, and HPs, with bound parameters determined by their operational characteristics, as detailed in [10]. The formulation (1) aligns with the flexibility model (2)-(4) since it calculates the flexibility cost based on the activated adjustment ranges in both power and accumulated energy consumption. By properly setting the cost coefficient vectors $\hat{\mathbf{c}}_{\text{p}}$, $\check{\mathbf{c}}_{\text{p}}$, $\hat{\mathbf{c}}_{\text{e}}$, and $\check{\mathbf{c}}_{\text{e}}$, formula (1) can reflect the costs of DERs offering flexibility. The next subsection will discuss the specification of these coefficients for typical resources.
@@ II-B Specification of the Cost Coefficients for Typical DERs
For resources whose flexibility can be described by the power adjustment range in each time slot (such as distributed PV and curtailable loads), flexibility cost can also be quantified using power adjustment ranges in each time slot. Therefore, the specification of the cost coefficients becomes straightforward: set $\hat{\mathbf{c}}_{\text{e}}$ and $\check{\mathbf{c}}_{\text{e}}$ to zero and specify $\hat{\mathbf{c}}_{\text{p}}$ and $\check{\mathbf{c}}_{\text{p}}$ using the same manner as traditional power-range-based formulations.
For resources with batteries such as EVs and BESSs, providing flexibility may make their charging and discharging process repeated or discontinuous, leading to battery aging. Given the energy trajectory, the battery aging can be estimated by the rainflow counting method [21, 22]. Since an EV or BESS providing flexibility means that its energy trajectory can be adjusted within the energy adjustment range, the device degradation part of its flexibility cost can be estimated by counting the expected battery aging cost on typical control signals, and the values of $\hat{\mathbf{c}}_{\text{e}}$ and $\check{\mathbf{c}}_{\text{e}}$ can be set accordingly.
In addition to the equipment degradation costs, flexibility costs also include the inconvenience caused to the DER users by the operational adjustment. This aspect can be directly perceived by DER users when they offer flexibility to the power system, yet it has not been adequately discussed in the literature. Therefore, the subsequent part of this subsection focuses on how to specify the flexibility cost coefficients for storage-like DERs, including EVs, BESS, and HPs, based on the potential inconvenience caused to their users.
@@ II-B1 Electric Vehicles
The flexibility of EVs lies in the charging process during their plug-in periods. Since the primary goal of EV charging is to charge the battery to the expected energy level by departure, the inconvenience caused to EV users by operational adjustment is reflected in the potentially unsatisfied charging demand at the departure time.
Fig. 2 illustrates a typical EV’s adjustment ranges in the accumulated energy consumption and its flexibility cost coefficients. In this example, the EV has an expected value for the energy to be charged by departure but also allows it to be adjustable within certain upper and lower bounds (the green and red dashed lines). The upper bound typically corresponds to the battery’s capacity limit, while the lower bound represents the minimum energy level acceptable to the EV user. The detailed calculation process of the upper and lower energy bound trajectories of EVs, as well as the subsequent BESSs and HPs, can be found in [10] and is omitted in this paper. The baseline energy trajectory is the energy trajectory where charging starts immediately at the rated power upon plug-in until the expected energy level is reached. The area within the upward and downward energy adjustment ranges represents the activated flexibility in accumulated energy consumption, which is where the actual energy trajectory can be adjusted. However, only the energy adjustment at the departure time can cause inconvenience to the EV user. More precisely, it is the downward adjustment that matters, as the upward adjustment, representing charging more than the expected energy, typically does not lead to an inconvenience in the EV’s usage. Thus, in the flexibility cost coefficients, only the coefficient for the downward energy adjustment range $\check{\mathbf{c}}_{\text{e}}$ at the departure time is non-zero, while $\check{\mathbf{c}}_{\text{e}}$ at other times, $\hat{\mathbf{c}}_{\text{e}}$, $\check{\mathbf{c}}_{\text{p}}$ and $\hat{\mathbf{c}}_{\text{p}}$ are all zero. [^3] The increased energy cost due to the overcharging indicated by the upward adjustments in the energy trajectory is not considered within the scope of flexibility costs. Instead, these costs will be incorporated into the energy cost component in the objective function of the flexibility activation program.
@@
In summary, based on the flexibility cost model (1), the aggregator can establish a contract with EV users that specifies compensation for any unsatisfied charging demands at the departure time. This contract directly correlates to the value of $\check{\mathbf{c}}_{\text{e}}$. Consequently, EV users can intuitively understand the potential inconvenience to their EV usage due to operational adjustment and assess their expected benefits. This level of clarity and directness is not achievable with traditional cost quantification methods based on power adjustment ranges, as power adjustments during charging are invisible to EV users.